A 20.00-ml sample of aqueous oxalic acid, H.C.O., was titrated with a 0.09113-M solution of potassium permanganate, KMnO..


2MnO4 (ag) + 5H₂C₂O₁(ag) + 6H* (ag) 1000₂(g) + 2Mn²+ (aq) + 8H₂O



A volume of 23.24 ml was required to reach the end point. What is the molarity of axalic acid ?


1
Expert's answer
2021-04-25T23:10:48-0400

Molar Mass of potassium permanganate = 158.034

= 0.09113 × 158.034 = 14.4

=20 + 14.4 + 23.24

= 57.642g

Molarity = moles of solute/Litres of the Solution

= 0.09113/0.0576

= 1.582m/L